# StoneDetector: Conventional and versatile code clone detection for Java

### v1

@@ 4.1 Benchmarks
A benchmark including in particular harder-to-detect code clones has been defined with GPTCloneBench [16]. The authors of the benchmark used the code clones of the SemanticBench [46] benchmark in prompts for generating additional code clones with the GPT-3 large language model. Manually validating the generated code clones afterward resulted in a set of 37,14 code clones, of which 16,989 are clone pairs with Java source code that we can use in our experiments. In spite of the benchmark’s authors intentions, we use the 5,431 near-miss code clones of Type 1 and Type 2 as well as the 11,558 Type 3 and Type 4 code clones in the experiments. Google Code Jam was an international programming contest where contestents were solving programming tasks. The posted solutions are publicly available, comprising 3.5 million code fragments in different programming languages for 716 programming tasks. Two different solutions to the same programming task can apparently be considered to share the same computations and thus to constitute a code clone. The thus defined code clones can vary across the different types of code clones. We consider solutions for the year 2022, and, as only part of the solutions are Java code, we are able thus to use 10,936,986 code clone pairs in our experiments. The last benchmark also originates from programming contests. Project CodeNet [15] assembles 14 million code fragments and more than 500 million lines of code with solutions to 4053 programming tasks, written in 55 programming languages. Different solutions to the same programming task are again considered code clones, however, near-miss code clones, i.e., code fragments sharing above 90% syntactical similarity, are omitted in this benchmark [15]. Around 5% of the code fragments, or 25 million lines of code, in Project CodeNet are Java code, which results in 11,212,500 code clone pairs used in our experiments. Note that all three aforementioned benchmarks, in contrast to BigCloneBench, do not further distinguish between the different clone types.
@@
Deckard SourcererCC NiCad iClones CCAligner StoneDetector NIL Oreo
Recall 18,318 20,827 20,829 20,842 20,831 20,828 20,844 20,844
T 1 87.8% 99.9% 99.9% 99.9% 99.9% 99.9% 100% 100%
Recall 3,317 3,444 3,460 3,453 3,456 3,475 3,460 3,475
T 2 95.4% 99.1% 99.5% 99.3% 99.4% 99.9% 99.5% 99.9%
Recall 2,139 3,209 3,243 3,253 3,218 3,209 3,271 3,283
VST 3 65.0% 97.6% 98.6% 98.9% 97.8% 97.6% 99.5% 99.8%
Recall 1,806 5,784 6,246 6,531 5,943 6,808 6,549 6,584
ST 3 24.4% 78.1% 84.3% 88.2% 80.2% 92.0% 88.5% 88.9%
Recall 2,311 1,462 169 4,051 1,784 6,324 3,499 5,187
MT 3 8.6% 5.5% 0.6% 15.1% 6.6% 23.6% 13.1% 19.4%
Recall 57,063 250 7 6,554 2,133 6,061 3,821 6,821
T 4 1.3% 0.0% 0.0% 0.1% 0.0% 0.1% 0.1% 0.1%
#Clones 13,239,074 804,315 754,022 987,659 1,244,160 1,300,458 1,274,852 1,015,425
Precision 46% 100% 100% 97% 98% 99% 99% 99%
@@
4.2 Research Question 1: Recall
@@
In order to address the first research question about how many code clones in Java source code our clone detection method is able to find, we ran StoneDetector on BigCloneBench via BigCloneEval [13]. For a comparison with the state of the art, we also ran seven other clone detectors on the same benchmark. The resulting numbers of identified known code clones, i.e., recall values, for the tools Deckard, SourcererCC, NiCad, iClones, CCAligner, StoneDetector, NIL, and Oreo are shown in Table 4, as measured by BigCloneEval. Since these numbers reflect only the annotated code clones in the benchmark (cf. also Sect. 4.1), we additionally list the overall number of detected clones. As noted before, we only consider clones with a minimum size of 15 lines of code. Duplicate clones, as reported by some of the tools, have been removed. Recall values are given in absolute and relative numbers.
@@
As can be seen in the table, all the code clone detectors, with the exception of Deckard, perform well on exact and near-miss code clones, i.e., code clones of Type 1, Type 2, and Very-Strongly Type 3. The respective recall values are located in the interval 97-100% and only for Deckard do we observe a worse recall value of 65%. These numbers in general reflect the results of prior research [18]. Thus, we can state that StoneDetector performs competitively with the state of the art for exact and near-miss code clones.
@@
As expected, beginning with code clones of Strongly Type 3, recall values are declining due to the larger syntactical differences of code fragments. We can also observe a difference between near-miss code clone detectors, e.g., NiCad and SourcererCC, and tools addressing harder-to-find clones, i.e., CCAligner, iClones, Oreo, StoneDetector, and NIL. StoneDetector achieves the best recall values for code clones of Strongly Type 3 and Moderately Type 3, but is overtaken by the tools Deckard, iClones, and Oreo for code clones of Type 4. However, the high number of Type 4 clones Deckard is finding has to be taken with care, considering its poor precision as described in the following section and also reported in the literature [18, 10].
When compared with state-of-the-art conventional clone detectors for Java source code, StoneDetector detects code clones with competitive recall and finds more code clones with 30% to 50% syntactical variance.
4.3 Research Question 2: Precision
@@
In addition to a high number of identified code clones, a code clone detector shall also avoid reporting on wrong code clones. Thus, we have to measure the clone detector’s precision, i.e., how many of the reported code clones are actual code clones. Unfortunately, answering the second research question is challenging, since BigCloneBench/BigCloneEval only contains information about a subset of all possible code clones in the benchmark. Due to the benchmark’s construction, only code clones of the considered 43 functionalities are known [12]. We have therefore performed a manual evaluation on random samples of the tool’s reported code clones, as is also usually conducted in the respective research literature [10, 11, 18].
@@
To this end, we randomly selected samples of 400 clone pairs of the overall reported code clones for each clone detector. The clone pairs were blindly, i.e., without revealing the respective tool, given to the three authors of this paper, who acted as judges. Each judge individually assessed the given clone pairs as actual or wrong clone pairs (true/false positive). Ambiguous assessments between the three judges were discussed and resolved afterward.
@@
StoneDetector detects code clones in Java source code, similar to other state-of-the-art clone detectors, at very high precision.
@@
We conducted further studies on the characteristics of the code clones found by StoneDetector and the other clone detectors. To this end, we made pairwise comparisons of the detected code clones for StoneDetector and each other tool, respectively, and looked into their common and exclusively identified code clones. On the one hand, we were able to observe a large fraction of commonly identified code clones. This mostly points to the subset of exact and near-miss clones as identified by all clone detectors (cf. also Table 4). On the other hand, we also found a significant portion of code clones with larger syntactical variance, in particular structural code clones and subclones, that are exclusively identified by StoneDetector and not by the other tools. For more details, we refer the reader to our previous study in [21].
4.4 Research Question 3: Generalizability
@@
Our third research question addresses the capability of clone detectors to identify code clones in source code with larger syntactical variability, in particular code clones of Strongly Type 3 and beyond. Besides, we are also interested in analyzing the clone detectors’ performance on benchmarks other than BigCloneBench in order to prevent the evaluation from being biased toward this particular benchmark. We therefore ran the clone detectors SourcererCC, NiCad, CCAligner, StoneDetector, NIL, and Oreo on the three benchmarks GPTCloneBench, Google Code Jam, and Project CodeNet, respectively, and report the observed recall values in Table 5. We omitted Deckard and iClones due to their lower precision compared to the other tools. We again limit clones to those with a minimal size of 15 lines, use the same configurations, and depict recall in absolute and relative numbers.
Benchmark SourcererCC NiCad CCAligner StoneDetector NIL Oreo
GPTCloneBench 4,577 3,646 3,452 4,068 4,681 4,366
26.9% 21.5% 20.3% 23.9% 27.6% 25.7%
Google Code Jam 101,843 54,363 102,129 362,989 212,597 115,879
0.9% 0.5% 0.9% 3.3% 1.9% 1.1%
Project CodeNet 104,790 121,271 145,218 713,942 646,232 537,322
0.9% 1.1% 1.4% 6.4% 5.8% 4.8%
@@
As shown in the table, recall values are in general at the lower end of the spectrum, which is, however, expected due to the nature of the benchmarks (cf. Sect. 4.1). In particular, the benchmarks Google Code Jam and Project CodeNet contain code clones, which originate from heterologously developed code and therefore usually include semantic code clones with differing syntax but similar computations. Thus, the recall values for all tools do not exceed 10% for both benchmarks. We can as well observe a distinction between the near-miss clone detectors NiCad and SourcererCC and tools addressing more difficult code clones, i.e., CCAligner, NIL, Oreo, and StoneDetector, where the latter ones perform better. While the contrast is not so apparent on the benchmark GPTCloneBench, due to its inclusion of Type 1/2 clones, it becomes obvious for the other two benchmarks. We also observe StoneDetector to perform the best on these two benchmarks.
@@
Tool 1 million LOC 10 million LOC 100 million LOC
NiCad 57s 18.49min 2.45h
CCAligner 18s 6.58min -
SourcererCC 4.17min 28.09min 46.52h
iClones 24s 7.26min -
StoneDetector 24s 6.17min 7.13h
Oreo 2.42min 9.54min 7.10h
@@
4.5 Research Question 4: Scalability

### v2

@@ 4.1 Benchmarks
The usage of BigCloneBench as dataset for harder-to-detect code clones of Type 3 and Type 4 have been controversially discussed in recent years[58, 53]. In particular, its naive usage as dataset for Deep learning clone detectors can compromise training and evaluation [92, 52]. The benchmark suffers from several issues, which are caused by its construction. For example, the benchmark is not only skewed towards Type 4 code clones but also towards certain functionalities, where, e.g., the Copy file functionality holds half of all code clones [53]. Besides, its ground truth quality, i.e., true code clones labels, is disputed for Type 4 code clones [58, 53]. However, note that these issues do not impair its usage for evaluating clone detectors on Type-1, Type-2, and Type-3 clones. However, in order to underpin our findings for harder-to-detect code clones of Type 3 and Type 4, we augment our evaluation with other state-of-the-art benchmarks for code clone detection.
@@
A benchmark including in particular harder-to-detect code clones has been defined with GPTCloneBench [4]. The authors of the benchmark used the code clones of the SemanticBench [3] benchmark in prompts for generating additional code clones with the GPT-3 large language model. Manually validating the generated code clones afterward resulted in a set of 37,149 code clones, of which 16,989 are clone pairs with Java source code that we can use in our experiments. In spite of the benchmark’s authors intentions, we use the 5,431 near-miss code clones of Type 1 and Type 2 as well as the 11,558 Type 3 and Type 4 code clones in the experiments. Google Code Jam was an international programming contest where contestents were solving programming tasks. The posted solutions are publicly available, comprising 3.5 million code fragments in different programming languages for 716 programming tasks. Two different solutions to the same programming task can apparently be considered to share the same computations and thus to constitute a code clone. The thus defined code clones can vary across the different types of code clones. We consider solutions for the year 2022, and, as only part of the solutions are Java code, we are able thus to use 10,936,986 code clone pairs in our experiments. Note that Google Code Jam also suffers from being biased towards certain dominant functionalities, i.e., programming tasks [58]. The last benchmark also originates from programming contests. Project CodeNet [78] assembles 14 million code fragments and more than 500 million lines of code with solutions to 4053 programming tasks, written in 55 programming languages. Different solutions to the same programming task are again considered code clones, however, near-miss code clones, i.e., code fragments sharing above 90% syntactical similarity, are omitted in this benchmark [78]. Around 5% of the code fragments, or 25 million lines of code, in Project CodeNet are Java code, which results in 11,212,500 code clone pairs used in our experiments. Apparently, in case of Google Code Jam and Project CodeNet, code clone samples from programming contests may not be representative of production code [58]. We though believe that using a wide range of benchmarks allows for a generalizable interpretation of our evaluation results, even if we are aware of the benchmarks’ issues. Note that all three benchmarks, in contrast to BigCloneBench, do not distinguish between the different clone types.
@@
Deckard SourcererCC NiCad iClones CCAligner StoneDetector NIL CloneWorks Oreo
Recall 18,318 20,827 20,829 20,842 20,831 20,828 20,844 20,828 20,844
T 1 87.8% 99.9% 99.9% 99.9% 99.9% 99.9% 100% 99.9% 100%
Recall 3,317 3,444 3,460 3,453 3,456 3,475 3,460 3,460 3,475
T 2 95.4% 99.1% 99.5% 99.3% 99.4% 99.9% 99.5% 99.5% 99.9%
Recall 2,139 3,209 3,243 3,253 3,218 3,209 3,271 3,240 3,283
VST 3 65.0% 97.6% 98.6% 98.9% 97.8% 97.6% 99.5% 98.5% 99.8%
Recall 1,806 5,784 6,246 6,531 5,943 6,808 6,549 7,073 6,584
ST 3 24.4% 78.1% 84.3% 88.2% 80.2% 92.0% 88.5% 95.5% 88.9%
Recall 2,311 1,462 169 4,051 1,784 6,324 3,499 1,305 5,187
MT 3 8.6% 5.5% 0.6% 15.1% 6.6% 23.6% 13.1% 4.8% 19.4%
Recall 57,063 250 7 6,554 2,133 6,061 3,821 19 6,821
T 4 1.3% 0.0% 0.0% 0.1% 0.0% 0.1% 0.1% 0.0% 0.1%
#Clones 13,239,074 804,315 754,022 987,659 1,244,160 1,300,458 1,274,852 915,434 1,015,425
Precision 46% 100% 100% 97% 98% 99% 99% 100% 99%
@@
4.2 Research Question 1: Recall and Precision
@@
In order to address the first research question about how many code clones in Java source code our clone detection method is able to find, we ran StoneDetector on BigCloneBench via BigCloneEval [103]. For a comparison with the state of the art, we also ran seven other clone detectors on the same benchmark. The resulting number of identified known code clones, i.e., recall values, for the tools Deckard, SourcererCC, NiCad, iClones, CCAligner, StoneDetector, NIL, CloneWorks, and Oreo are shown in Table 4, as measured by BigCloneEval. Since these numbers reflect only the annotated code clones in the benchmark (cf. Sect. 4.1), we additionally list the overall number of detected clones. As noted before, we only consider clones with a minimum size of 15 lines of code. Duplicate clones, as reported by some tools, were removed. Recall values are given in absolute and relative numbers.
@@
As can be seen in the table, all the code clone detectors, with the exception of Deckard, perform well on exact and near-miss code clones, i.e., code clones of Type 1, Type 2, and Very-Strongly Type 3. The respective recall values are located in the interval 97-100% and only for Deckard do we observe a worse recall value of 65%. These numbers in general reflect the results of prior research [88]. Thus, we can state that StoneDetector performs competitively with the state of the art for exact and near-miss code clones. Note though that in case of Type 1 clones, i.e., exact duplicates of code fragments that only differ in layout, comments, or formatting (cf. Sect 1), StoneDetector misses certain code clones. Closer inspection revealed that these cases can be attributed to a low number of code samples from the benchmark which can not be properly processed or parsed using the Spoon analysis framework.
@@
As expected, beginning with code clones of Strongly Type 3, recall values are declining due to the larger syntactical differences of code fragments. We can also observe a difference between near-miss code clone detectors, e.g., NiCad, CloneWorks, and SourcererCC, and tools addressing harder-to-find clones, i.e., CCAligner, iClones, Oreo, StoneDetector, and NIL. StoneDetector achieves the best recall values for code clones of Strongly Type 3 and Moderately Type 3, but is overtaken by the tools Deckard, iClones, and Oreo for code clones of Type 4. However, the high number of Type 4 clones Deckard is finding has to be taken with care, considering its poor precision as described in the following section and also reported in the literature [88, 110].
@@
In addition to a high number of identified code clones, a code clone detector shall also avoid reporting on wrong code clones. Thus, we have to measure the clone detector’s precision, i.e., how many of the reported code clones are actual code clones. Unfortunately, answering this question is challenging, since BigCloneBench/BigCloneEval only contains information about a subset of all possible code clones in the benchmark. Due to the benchmark’s construction, only code clones of the considered 43 functionalities are known [102]. We have therefore performed a manual evaluation on random samples of the tool’s reported code clones, as is also usually conducted in the respective research literature [110, 87, 88].
@@
To this end, we randomly selected samples of 400 clone pairs of the overall code reported clones for each clone detector. The clone pairs were blindly, i.e., without revealing the respective tool, given to the three authors of this paper, who acted as judges. Each judge individually assessed the given clone pairs as actual or wrong clone pairs (true/false positive). Ambiguous assessments between the three judges were discussed and resolved afterward.
@@
When compared with state-of-the-art conventional clone detectors for Java source code, StoneDetector detects code clones with competitive recall and finds more code clones with 30% to 50% syntactical variance. At the same time, StoneDetector detects code clones in Java source code, similar to other state-of-the-art clone detectors, at very high precision.
4.2.1 Qualitative Analysis
@@
We conducted further studies on the characteristics of the code clones found by StoneDetector and the other clone detectors. To this end, we first made pairwise comparisons of the detected code clones for StoneDetector and some of the other tools, respectively, and looked into their common and exclusively identified code clones. On the one hand, we were thereby able to observe a large fraction of commonly identified code clones. This mostly points to the subset of exact and near-miss clones as identified by all clone detectors (cf. also Table 4). On the other hand, we also found a significant portion of code clones with larger syntactical variance, in particular structural code clones, i.e, code clones which implement the same control flow using different syntactical constructs, and subclones, i.e., code clones where the control flow of one code fragment is embedded into the other fragment’s control flow [8]. Code clones of these types are often exclusively identified by StoneDetector and not by the other tools.
When analyzing clone pairs that are detected by StoneDetector but not by the other tools, we created random samples of found code clones on BigCloneBench, specifically for comparing StoneDetector with NiCad, Oreo, CCAligner, and CloneWorks, as representatives of near-miss clone detectors, tools focusing on large-gap clones, and Deep Learning approaches. For this purpose, for each tool $A$, we generated a random sample of 100 clones for the set of code clones identified by StoneDetector but not by AA ($StoneDetector\setminus A$) and for the set of code clones detected by $A$ but not by StoneDetector ($A\setminus StoneDetector$), respectively, and manually inspected those samples. Finally, we also created a random sample of 100 code clones which were only detected by StoneDetector and not by any other tool considered in our evaluation, though excluding Deckard due to its low precision (i.e., NiCad, SourcererCC, CCAligner, iClones, NIL, Oreo, and CloneWorks).
StoneDetector vs NiCad
The closer examination of 100 random clone pairs that were detected by StoneDetector but not by NiCad revealed that StoneDetector finds more clones with greater syntactic variance and in particular, structural clones (55) and subclones (35). The remaining clones were classified as of other types (4) and false positives (6). Analysis of 100 clones found only by NiCad revealed that these were mainly due to NiCad’s pretty printing, which often leads to an increase in code size and thus higher permissiveness for code clones, and due to different path sizes in the dominator trees in StoneDetector and token sequences in NiCad.
StoneDetector vs CloneWorks
In CloneWorks, clone detection is based on a modified Jaccard metric. This metric represents code fragments as sets of contained tokens and calculates their minimum rate of overlapping tokens [104]. Similar to Oreo, CloneWorks thus does not relies on the order of code fragment’s instructions for clone detection. As a result, CloneWorks can also detect simple structural clones and subclones, if the lengths of the code fragments do not differ too much. Our examination of clone pairs found exclusively by StoneDetector showed that more structural clones (28) and subclones (64) were detected. A closer inspection revealed that the clones often differed greatly in the number of lines of code. Further clones found exclusively by StoneDetector were classified as of other clone types (6) and false positives (2). Clone pairs found exclusively by CloneWorks can be attributed to the use of token sets. Structural clones (24) and subclones (49) were consequently also found by CloneWorks if the order of the instructions in the clone pairs differs. StoneDetector cannot detect such clones because its used LCS method requires a sequence of instructions. For example, code fragments with a high number of identical instructions occurring at different code locations cause issues for LCS and are often not recognized as clones even though they contain the very same instructions. Other code clones found exclusively by CloneWorks were such clones, in which instructions appeared in short form in one code fragment and in full form in the other (7), e.g., as in case of a foreach loop and a normal for loop over iterators. In contrast to StoneDetector, which conceives the latter loop using four sub-expressions, CloneWorks treats both loops as one expression in its aggressive version. Remaining code clones found exclusively by CloneWorks were again of others clone types (17) and false positives (3).
StoneDetector vs CCAligner
Our manual inspection of code clones found by StoneDetector but not by CCAligner revealed that CCAligner misses many subclones, which are detected by StoneDetector (52), which can be attributed to its code window approach failing in case of scattered code modifications. Furthermore, StoneDetector finds structural code clones based on its applied control flow representation of dominator trees, which are missed by CCAligner (21). Remaining code clones were classified as of other clone types (18) and false positives (9). In contrast, CCAligner’s code window approach allowed for finding subclones with many concentrated code modifications which are missed by StoneDetector (26). CCAligner, in addition, considers a function inside an anonymous class and the surrounding code as potential code clones, which is not the case for StoneDetector (31). The remaining code clones exclusively detected by CCAligner were classified as of other types (33) and false positives(10).
StoneDetector vs Oreo
Oreo uses functional metrics for clone detection, which determine, for example, the number of field accesses, array accesses, loops, etc., in code snippets. Clone detection in Oreo does not depend on the order of the instructions, but instead analyzes code properties by means of these functional metrics. In the manual review of the random samples, we found many structural clones (49) and subclones (14) that were overlooked by Oreo but detected by StoneDetector. These were often code clones that were very different in structure, which StoneDetector recognizes, but for which Oreo derives metric measures that are too distant from each other. The remaining clones found exclusively by StoneDetector were classified as of other clone types (31) and false positives (6). On the other hand, most of the code clones found exclusively by Oreo were structural clones (52) that often contained subpaths in internal substructures which were disordered or of very different lengths. It was also noticeable that Oreo thereof identified subclones (25) consisting of basic blocks, although they have very different lengths or their instructions were of different order, which is not a problem for Oreo due to its usage of functional metrics but pose an issues for the string metric comparison in StoneDetector. Remaining clones found exclusively by Oreo were assigned as of other types (21) or as false positives (2).
StoneDetector vs All
Finally, we also looked into the set of code clones in BigCloneBench which were exclusively identified by StoneDetector but not by the other clone detectors considered in our evaluation. From the total of 1,300,458 code clones identified by StoneDetector (cf. Table 4), a fraction of 235,982 clones were not detected by any of the other tools, i.e., almost each fifth code clone (18%). A closer examination of 100 randomly selected clones of this set again showed that StoneDetector finds significantly more subclones (49) and structural clones (35), which often differed greatly in size. Additional clones were classified as other (14) and false (2).
4.3 Research Question 2: Generalizability
@@
Our second research question addresses the capability of clone detectors to identify code clones in source code with larger syntactical variability, in particular code clones of Strongly Type 3 and beyond. Besides, we are also interested in analyzing the clone detectors’ performance on benchmarks other than BigCloneBench in order to prevent the evaluation from being biased toward this particular benchmark, also considering the benchmark’s various issues as mentioned above. We therefore ran the clone detectors SourcererCC, NiCad, CCAligner, StoneDetector, NIL, CloneWorks, and Oreo on the three benchmarks GPTCloneBench, Google Code Jam, and Project CodeNet, respectively, and report the observed recall values in Table 5. We omitted Deckard and iClones due to their lower precision compared to the other tools. We again limit clones to those with a minimal size of 15 lines, use the same configurations, and depict recall in absolute and relative numbers.
Benchmark SourcererCC NiCad CCAligner StoneDetector NIL Oreo CloneWorks
GPTClone 4,577 3,646 3,452 4,068 4,681 4,366 3,987
Bench 26.9% 21.5% 20.3% 23.9% 27.6% 25.7% 23.4%
Google Code 101,843 54,363 102,129 362,989 212,597 115,879 91,794
Jam 0.9% 0.5% 0.9% 3.3% 1.9% 1.1% 0.8%
Project 104,790 121,271 145,218 713,942 646,232 537,322 266,888
CodeNet 0.9% 1.1% 1.4% 6.4% 5.8% 4.8% 2.4%
@@
As shown in the table, recall values are in general at the lower end of the spectrum, which is, however, expected due to the nature of the benchmarks (cf. Sect. 4.1). In particular, the benchmarks Google Code Jam and Project CodeNet contain code clones, which originate from heterologously developed code and therefore usually include semantic code clones with differing syntax but similar computations. Thus, the recall values for all tools do not exceed 10% for both benchmarks. We can as well observe a distinction between the near-miss clone detectors NiCad, CloneWorks, and SourcererCC and tools addressing more difficult code clones, i.e., CCAligner, NIL, Oreo, and StoneDetector, where the latter ones perform better. While the contrast is not so apparent on the benchmark GPTCloneBench, due to its inclusion of Type 1/2 clones, it becomes obvious for the other two benchmarks. We also observe StoneDetector to perform the best on these two benchmarks.
@@
Tool 1 million LOC 10 million LOC 100 million LOC 328 million LOC
NiCad 57s 18.49min 2.45h -
CCAligner 18s 6.58min - -
SourcererCC 4.17min 28.09min 46.52h -
iClones 24s 7.26min - -
StoneDetector 24s 6.17min 7.13h 79h
StoneDetector 24s 2.48min 1.11h 12.32h
Oreo 2.42min 9.54min 7.10h 23.53h
NIL 2s 17s 10.21min 1.44h
CloneWorks 10s 1.46min 25.35min 4.07h
@@
4.4 Research Question 3: Scalability